When a 4.348 g sample of NaOH(s) is dissolved in 100 g of water in a calorimeter, the temperature rises from 23.6ºC to 47.4ºC. Calculate HdissinkJ/mol.Assume that Ccalorimeter= 35.5J/0C.Remember, fordilute solutions, Csolnis Cwater= 4.18 J/g·0Ca.

a) Calculate the amount of heat absorbed by the solution(qsoln).

b)Calculate the amount of heat absorbed by the calorimeter(qcal).
Answer #1

NaOH moles used = mass / molar mass = 4.348 g / ( 40 g/mol) = 0.109 mol

Tempereature rise = 47.4 - 23.6 = 23.8 C

mass of solution = 4.348 + 100 = 104.348 g

a) Heat absorbed by solution qsoln = specific heat of solution or water x mass of solution x temp rise

= 4.18 J/gC x 104.348 g x 23.8 C

= 10381 J

b) heat absorbed by calorimeter = Ccalorimeter x temp rise = 35.5 J/C x 47.4C = 1683 J

Total heat absorbed = 10381 + 1683 = 12064 = 12.064 KJ

Heat of dissolution = heat absorbed by solution and calorimter / moles of NaOH

= - 12.064 KJ / 0.109 mol = - 111 KJ/mol

( -ve sign indicates heat released during dissolution)